# MHONGOOSE — A MeerKAT Nearby Galaxy H i Survey

### v1

@@ 6.2 Comparison with single dish profiles
We additionally list the HIPASS and GBT flux densities in Tab. 6 for those galaxies where at least one single-dish measurement is available. We have re-measured the GBT flux densities from the original profiles, as the values listed in Sardone et al. (2021) represent cumulative flux densities integrated over the entire respective areas mapped by the GBT, rather than direct integration of the global profiles as discussed here. The H i profiles themselves are shown and compared in Fig. 10. Where necessary we have shifted the literature spectra to the same radio velocity scale as the MeerKAT profiles. Any “peaks” detected in the single-dish data outside the velocity range of the target galaxies (e.g., the HIPASS profile of J0008–34 at $v\sim 270$ km s-1 in Fig. 10) are due to noise.
Figure 10: Comparison of the MeerKAT and single dish profiles. MeerKAT r15_t00 profiles are shown in black, GBT profiles (Sardone et al. 2021) in cyan, and HIPASS profiles (Koribalski et al. 2004) in red. The green arrows indicate the central velocties as derived from the MeerKAT data.
@@
The agreement with the re-measured GBT flux densities is good. The ratio of the GBT and MeerKAT flux densities is $1.058\pm 0.080$, so the GBT on average detects $\sim 6\%$ more flux, but this is within the scatter of $\sim 8\%$. A least-squares fit to the linear values, as shown in Fig.11, has a slope of 0.964, but this slope is mainly determined by NGC 7424 (J2257–41), the galaxy with the largest flux density, where we find that the GBT actually detects less H i than MeerKAT. Leaving that point out we end up with a slope of 0.997, again indicating good agreement between the MeerKAT and GBT fluxes. This is especially so when one considers the uncertainty in the MeerKAT flux calibration of $\sim 10\%$ (Serra et al. 2023) and a most likely similar uncertainty in the GBT flux calibration. The good agreement is consistent with the fact that most sample galaxies are significantly less extended than the $\sim 20^{\prime}$ maximum angular scale that the MeerKAT shortest baseline of 29m is sensitive to.
@@ 7.1 Pixel-pixel density histograms of individual galaxies
An efficient method to further summarise this information is pixel-by-pixel scatter plots of the various moment values: for each location (pixel) the value of one moment is plotted against that of a different moment at that same location. A number of previous studies have used these plots to study properties of galaxies. For example, de Blok et al. (2018) used the distribution of the zeroth- and second-moment values of the H i distribution in the M81 triplet to find that each of the three triplet galaxies occupied unique positions in that particular parameter space, and addressed the question of whether the M81 triplet could be the equivalent of a Lyman-$\alpha$ absorber (cf. their Fig. 20). Maccagni et al. (2021) presented optical line emission data of Fornax A and used the distribution of zeroth- and first-moment values to distinguish between rotating and outflowing gas. Sun et al. (2018, 2020) studied the zeroth and second moments of the CO distribution in the Physics at High Angular resolution in Nearby GalaxieS (PHANGS) galaxies.
@@
Figure 12: Two-dimensional density histograms showing the distribution of inclination-corrected column densities and second moment values for the MHONGOOSE galaxies. Each histogram consists of $75\times 75$ hexagonal bins, and only bins containing more than 4 pixels are shown. The colour bar on the right in each panel indicates the number of pixels per bin. The background contours indicate the total distribution of the entire sample as shown in Fig. 13. Galaxies are divided in three bins in rotation velocity. Within each bin, galaxies are ordered by increasing inclination, with the most face-on galaxies in the top-left, and the most edge-on in the bottom-right. The two light-gray bars between the panels in each bin indicate inclinations of 50 $(b/a=0.67)$ and 70 $(b/a=0.33)$, respectively.
The panels in Fig. 12 show, for each galaxy, the number of pixels having values corresponding to any given combination of column density (horizontal axis) and velocity spread (vertical axis). This is shown as a two-dimensional density-histogram with each bin having the value given by the colour scale for that panel. Bins are only shown if there are a minimum of four moment pixels falling in them. The contours show a similar density histogram but in this case derived from all the single-track moment maps of all galaxies combined. The juxtaposition of the two distributions allows for a quick comparison of the distribution for a particular galaxy with that of the entire sample. This total distribution is further discussed in Sect. 7.2.
@@ 7.2 Total pixel-by-pixel density histogram
Figure 13: Two-dimensional density histograms showing the distribution of zeroth and second moment values for the entire MHONGOOSE sample. The histogram consists of $75\times 75$ hexagonal bins, and only bins containing more than 8 pixels are shown. The colour bar indicates the number of pixels with the H i column density and velocity widths indicated by the horizontal and vertical axes, respectively.
@@
At the highest column densities we see a steep drop-off in the density distribution, despite this being the part of the diagram with the highest signal-to-noise. This drop-off occurs at surface density values of $\log(N_{{\rm H{\hskip 0.56917pt\tt I}}}/{\rm cm}^{-2})\sim 21.1$, or $\sim 10$ $M_{\odot}\,{\rm pc}^{-2}$, and thus implies an upper limit to the H i column density (at this resolution). This value agrees well with the upper limit of $\approx 9$ $M_{\odot}\,{\rm pc}^{-2}$ found by Bigiel et al. (2008) for the THINGS sample at similar resolutions (also see Leroy et al. 2008; for a more detailed observational view, see Schruba et al. 2017). A comparison with molecular gas column density values shows that this upper limit corresponds with a transition from an atomic to a molecular ISM (cf. Krumholz et al. 2009; McKee & Krumholz 2010; Sternberg et al. 2014)
Turning to the second-moment values, we see that at surface densities $\log(N_{{\rm H{\hskip 0.56917pt\tt I}}}/{\rm cm}^{-2})>19.5$ there is a well-defined lower limit to the second-moment values. This limit occurs at $\sim 8.5$ km s-1. Given the 1.4 km s-1 velocity resolution of the data (corresponding to the lower limit of the plotted second-moment range) and the high signal-to-noise of the profiles in this part of the diagram, this cannot be a velocity resolution artefact but must indicate an absence of H i profiles with second-moment values less than $\sim 8.5$ km s-1. To first order, the ISM is considered to consist of a cool neutral medium (CNM) with a temperature of a few tens of K and a warm neutral medium (WNM) with temperatures around $\sim 8000$K . The CNM line profiles have a line width of $\sim 1$ km s-1 while WNM line width are $\sim 8$ km s-1 (in the absence of any other processes that can broaden line-widths) (Braun 1997; Wong & Blitz 2002; Wolfire et al. 2003). The lower limit found in our data does not indicate an absence of CNM, but merely shows that the H i emission, and therefore the line width, is dominated by the WNM, where the latter has a lower limit to its observed velocity dispersion as shown in Fig. 13.
@@
A possible explanation for the presence of this division is related to the formation of the CNM. This was put in a physical context in Kanekar et al. (2011). Based on measurements of the spin temperature in a large number of Galactic H i clouds they identify a rapid change in spin temperature at $\sim 2\cdot 10^{20}$ cm-2. They associate this with a threshold H i surface density for formation of the CNM. Above this $\sim 2\cdot 10^{20}$ cm-2 threshold, the H i can sufficiently self-shield against UV photons to enable formation of a cool component. Below this threshold the UV photons can heat and ionise the gas, preventing the formation of a CNM. Note that this threshold also corresponds to that used to define a Damped Lyman-$\alpha$ system.
@@ 7.3 Masking and moments
The SoFiA-2 moment map creation steps described above are a sophisticated combination of source finding and masking at different spatial and velocity scales. The essentials of this can, fortunately, already be captured by considering an idealised Gaussian H i profile with constant velocity dispersion and variable normalisation or peak value (corresponding to different column densities).
@@
We consider the second moment of the remaining $>4\sigma_{\rm ch}$ profile after masking as a function of the signal-to-noise of the profile. Without masking (and for noise-free data), this would equal the velocity dispersion of the Gaussian. With masking, it decreases as the peak value decreases.
@@
Note how well the down-turn in the tail is described by the simple single Gaussian model. We can deduce for the masking procedure described here, that at S/N values between $\sim 50$ and $\sim 10$, the second moment values start to underestimate the velocity dispersion. At a ${\rm S/N}\sim 5$, the second moment underestimates the velocity dispersion by a factor of $\sim 2.5$. Column densities, though not as affected, can nevertheless be underestimated by factors of $\sim 1.6$ at these lower S/N values. Multiple components along the line of sight will not significantly influence the lower bounds of the distribution as, at a given column density, these will be found at higher second-moment values. Nevertheless, fitting multiple Gaussians to the velocity profiles could further elucidate (or better: circumvent) the effects of masking in the presence of multiple profiles.
@@ 8 Other galaxies in the MHONGOOSE fields
The sensitivity and large field of view of MeerKAT result in the detection of additional companion and field galaxies. Here we give a first inventory of the galaxies detected in the 1.5 by 1.5 single-track fields in a velocity range of $-500$ to $+500$ km s-1 with respect to the central velocity of the target galaxies. Detections are based on the single-track zeroth-moment maps. A list of these additional galaxies is given in Table 7. For the fields where additional galaxies are present, Fig. 15 shows the positions of these galaxies. We expect an even larger number of galaxies to be detected in the full-depth data once available. While the search volume is defined with respect to our target galaxies, this does not imply that every one of these additional galaxies is also physically associated with those target galaxies. Most of them will be, but it is possible that (especially at the edges of the volume probed) we are picking up some foreground or background galaxies
@@
Three additional galaxies visible in the moment maps are partially outside of the volume of the data cubes used here. These are LEDA 583148 in the J0309–41 field, ESO 486–21 in the J0459–26 field, and MKT J125225.4-124304.1 in the J1253–12 field. For these three galaxies we list the position and approximate radial velocity in Table 7. We refrain from presenting their H i line widths and fluxes because the large uncertainty due to the incomplete coverage by the data used here. The last of these three galaxies is a fifth, previously uncatalogued galaxy. All of the additional galaxies listed here can thus be associated with an optical counterpart; we have not found any “dark” galaxies.
@@
A more complete census of the companion galaxy population will be possible when the full-depth data for the full sample are available and we can probe the companion population to the fullest extent possible for our survey. As with the galaxies presented here, these satellites and companions will be resolved spatially and spectrally, meaning that in-depth studies of the dark matter content and baryon fractions can be made down to an H i mass limit of $\sim 10^{6}$ $M_{\odot}$ (cf. Table 5), a regime previously only accessible in the Local Group. A first discovery of such a low-mass galaxy in the MHONGOOSE observations is presented in Maccagni et al. (subm.).
@@
 [^11] Galaxies in a bold font match their optical counterparts both in position and in velocity. Galaxies in a normal font match the catalogued optical counterpart in position, but with no prior catalogued redshift. For these galaxies our data give the first redshift measurements. For the galaxies in an italic font, there is a spatial match with a previously uncatalogued optical counterpart. The “Galaxies with uncertain parameters” are only partially included in the data presented here. The central velocities listed are approximate.
@@ 9 Summary
We have presented a description of the MHONGOOSE survey, a large, deep, high-resolution H i survey on MeerKAT of 30 nearby gas-rich spiral and dwarf galaxies. The total survey comprises 1650h of observing time, with 55h allocated per galaxy. The main science goal of MHONGOOSE is to characterise the low-column density H i in these galaxies, linking it with various cold gas accretion scenarios. Other science goals include probing in detail the link between H i and star formation, the kinematics of the galaxies, constraining the presence and distribution of dark matter, as well as the environments of the targets. The sample covers $\sim 3$ orders of magnitude in H i mass, from low-mass $M_{\rm H{\hskip 0.56917pt\tt I}}\sim 10^{7}$ $M_{\odot}$ as typically found in dwarf galaxies to $M_{\rm H{\hskip 0.56917pt\tt I}}\sim 10^{10}$ $M_{\odot}$ as found in large spiral galaxies.
In this paper, we present the survey, the sample, and a summary of the observations. In addition we describe the data reduction procedure and data products.
@@
Each galaxy will eventually be observed for 55h in 10 tracks of 5.5h, and we produce full-depth data cubes at a velocity resolution of 1.4 km s-1 and a range in angular resolutions from $\sim 8^{\prime\prime}$ as the highest resolution to $\sim 90^{\prime\prime}$ as our lowest resolution. The resulting column density sensitivity ($3\sigma$ over 16 km s-1) ranges from $4.3\cdot 10^{19}$ cm-2 at the highest resolution to $4.2\cdot 10^{17}$ cm-2 at the lowest resolution. The $3\sigma$, 50 km s-1 mass-detection limit is $5.4\cdot 10^{5}$ $M_{\odot}$ at 10 Mpc, the median distance of the sample.
@@
We present first results based on the full-depth data of a limited number of galaxies and single-track observations of the complete sample (the latter observations comprising 10% of the survey data). Comparing full-depth, single-track and single-dish data, we show that MeerKAT is excellent at recovering the total fluxes of our sample galaxies. We also show that the full-depth data detect on average $\sim 2\%$ more H i than the single-track data. While this increase may seem modest, the additional H i mass it represents is sufficient to allow for the existence of a $\sim 10^{18}$ cm-2 low-column density H i component in the disks of our galaxies. This shows the difficulty of investigating the presence of low-column density H i using total H i mass measurements.
@@
We also investigate the link between the local H i surface density and the velocity “dispersion” in all of our sample galaxies. We recover the well-known upper limit on H i surface density (measured on kpc scales), and find a lower limit to the measured second-moment value of $\sim 8.5$ km s-1. We find that the H i surface density has two distinct locations in a surface density-velocity dispersion diagram. One at $\sim 5\cdot 10^{20}$ cm-2, representing the gas in the star-forming disks, and one at $\sim 5\cdot 10^{19}$ cm-2 representing gas not participating in star formation. The two regions are distinct and separated at an H i surface density of $2\cdot 10^{20}$ cm-2. Previous studies (e.g. Kanekar et al. 2011) have shown that at this surface density a rapid change in spin temperature occurs, and that this is the column density where the formation of a cool and dense component of the ISM starts.
@@
We give an overview of all additional galaxies detected in a the single-track data. These are defined as the sources detected in H i that were not targeted by the survey, but were nevertheless found within the $1.5^{\circ}\times 1.5^{\circ}$ area and spanning 1000 km s-1 in velocity range around the target galaxies. We find that half of these 49 galaxies have optical and catalogued counterparts that coincide in position and velocity. Just under half have catalogued optical counterparts that coincide in position but that had until now no previously measured redshifts. Five of our additional galaxies have optical counterparts that have not been previously catalogued. We have not found any H i detections that do not have optical counterparts.
@@
Studies of the full-depth data will be presented in future papers. However, based on the results shown here, it is already clear that the MHONGOOSE data is of excellent quality, showing great promise for future work on the low-column density H i in nearby galaxies.

### v2

@@ 6.2 Comparison with single-dish profiles
We additionally list the HIPASS and GBT flux densities in Tab. 6 for those galaxies where at least one single-dish measurement is available. We have remeasured the GBT flux densities from the original profiles, as the values listed in Sardone et al. (2021) represent cumulative flux densities integrated over the entire respective areas mapped by the GBT, rather than direct integration of the global profiles as discussed here. The H i profiles themselves are shown and compared in Fig. 10. Where necessary we have shifted the literature spectra to the same radio velocity scale as the MeerKAT profiles. Any “peaks” detected in the single-dish data outside the velocity range of the target galaxies (e.g., the HIPASS profile of J0008–34 at $v\sim 270$ km s-1 in Fig. 10) are due to noise.
Figure 10: Comparison of the MeerKAT and single-dish profiles. MeerKAT r15_t00 profiles are shown in black, GBT profiles (Sardone et al. 2021) in cyan, and HIPASS profiles (Koribalski et al. 2004) in red. The green arrows indicate the central velocties as derived from the MeerKAT data.
@@
The agreement with the remeasured GBT flux densities is good. The ratio of the GBT and MeerKAT flux densities is $1.058\pm 0.080$, so the GBT on average detects $\sim 6\%$ more flux, but this is within the scatter of $\sim 8\%$. A least-squares fit to the linear values, as shown in Fig.11, has a slope of 0.964, but this slope is mainly determined by NGC 7424 (J2257–41), the galaxy with the largest flux density, where we find that the GBT actually detects less H i than MeerKAT. Leaving that point out we end up with a slope of 0.997, again indicating good agreement between the MeerKAT and GBT fluxes. This is especially so when one considers the uncertainty in the MeerKAT flux calibration of $\sim 10\%$ (Serra et al. 2023) and a most likely similar uncertainty in the GBT flux calibration. The good agreement is consistent with the fact that most sample galaxies are significantly less extended than the $\sim 20^{\prime}$ maximum angular scale that the MeerKAT shortest baseline of 29m is sensitive to.
@@ 7.1 Pixel-pixel density histograms of individual galaxies
An efficient method to further summarize this information is pixel-by-pixel scatter plots of the various moment values: for each location (pixel) the value of one moment is plotted against that of a different moment at that same location. A number of previous studies have used these plots to study properties of galaxies. For example, de Blok et al. (2018) used the distribution of the zeroth- and second-moment values of the H i distribution in the M81 triplet to find that each of the three triplet galaxies occupied unique positions in that particular parameter space, and addressed the question of whether the M81 triplet could be the equivalent of a Lyman-$\alpha$ absorber (cf. their Fig. 20). Maccagni et al. (2021) presented optical line emission data of Fornax A and used the distribution of zeroth- and first-moment values to distinguish between rotating and outflowing gas. Sun et al. (2018, 2020) studied the zeroth and second moments of the CO distribution in the Physics at High Angular resolution in Nearby GalaxieS (PHANGS) galaxies.
@@
Figure 12: Two-dimensional density histograms showing the distribution of inclination-corrected column densities and second moment values for the MHONGOOSE galaxies. Each histogram consists of $75\times 75$ hexagonal bins, and only bins containing more than 4 pixels are shown. The color bar on the right in each panel indicates the number of pixels per bin. The background contours indicate the total distribution of the entire sample as shown in Fig. 13. Galaxies are divided in three bins in rotation velocity. Within each bin, galaxies are ordered by increasing inclination, with the most face-on galaxies in the top-left, and the most edge-on in the bottom-right. The two light-gray bars between the panels in each bin indicate inclinations of 50 $(b/a=0.67)$ and 70 $(b/a=0.33)$, respectively.
The panels in Fig. 12 show, for each galaxy, the number of pixels having values corresponding to any given combination of column density (horizontal axis) and velocity spread (vertical axis). This is shown as a two-dimensional density-histogram with each bin having the value given by the color scale for that panel. Bins are only shown if there are a minimum of four moment pixels falling in them. The contours show a similar density histogram but in this case derived from all the single-track moment maps of all galaxies combined. The juxtaposition of the two distributions allows for a quick comparison of the distribution for a particular galaxy with that of the entire sample. This total distribution is further discussed in Sect. 7.2.
@@ 7.2 Total pixel-by-pixel density histogram
Figure 13: Two-dimensional density histograms showing the distribution of zeroth and second moment values for the entire MHONGOOSE sample. The histogram consists of $75\times 75$ hexagonal bins, and only bins containing more than 8 pixels are shown. The color bar indicates the number of pixels with the H i column density and velocity widths indicated by the horizontal and vertical axes, respectively.
@@
At the highest column densities we see a steep drop-off in the density distribution, despite this being the part of the diagram with the highest S/N. This drop-off occurs at surface density values of $\log(N_{{\rm H{\hskip 0.56917pt\tt I}}}/{\rm cm}^{-2})\sim 21.1$, or $\sim 10$ $M_{\odot}\,{\rm pc}^{-2}$, and thus implies an upper limit to the H i column density (at this resolution). This value agrees well with the upper limit of $\approx 9$ $M_{\odot}\,{\rm pc}^{-2}$ found by Bigiel et al. (2008) for the THINGS sample at similar resolutions (also see Leroy et al. 2008; for a more detailed observational view, see Schruba et al. 2017). A comparison with molecular gas column density values shows that this upper limit corresponds with a transition from an atomic to a molecular ISM (cf. Krumholz et al. 2009; McKee & Krumholz 2010; Sternberg et al. 2014)
Turning to the second-moment values, we see that at surface densities $\log(N_{{\rm H{\hskip 0.56917pt\tt I}}}/{\rm cm}^{-2})>19.5$ there is a well-defined lower limit to the second-moment values. This limit occurs at $\sim 8.5$ km s-1. Given the 1.4 km s-1 velocity resolution of the data (corresponding to the lower limit of the plotted second-moment range) and the high S/N of the profiles in this part of the diagram, this cannot be a velocity resolution artefact but must indicate an absence of H i profiles with second-moment values less than $\sim 8.5$ km s-1. To first order, the ISM is considered to consist of a cool neutral medium (CNM) with a temperature of a few tens of K and a warm neutral medium (WNM) with temperatures around $\sim 8000$K . The CNM line profiles have a line width of $\sim 1$ km s-1 while WNM line width are $\sim 8$ km s-1 (in the absence of any other processes that can broaden line-widths) (Braun 1997; Wong & Blitz 2002; Wolfire et al. 2003). The lower limit found in our data does not indicate an absence of CNM, but merely shows that the H i emission, and therefore the line width, is dominated by the WNM, where the latter has a lower limit to its observed velocity dispersion as shown in Fig. 13.
@@
A possible explanation for the presence of this division is related to the formation of the CNM. This was put in a physical context in Kanekar et al. (2011). Based on measurements of the spin temperature in a large number of Galactic H i clouds they identify a rapid change in spin temperature at $\sim 2\cdot 10^{20}$ cm-2. They associate this with a threshold H i surface density for formation of the CNM. Above this $\sim 2\cdot 10^{20}$ cm-2 threshold, the H i can sufficiently self-shield against UV photons to enable formation of a cool component. Below this threshold the UV photons can heat and ionize the gas, preventing the formation of a CNM. Note that this threshold also corresponds to that used to define a Damped Lyman-$\alpha$ system.
@@ 7.3 Masking and moments
The SoFiA-2 moment map creation steps described above are a sophisticated combination of source finding and masking at different spatial and velocity scales. The essentials of this can, fortunately, already be captured by considering an idealized Gaussian H i profile with constant velocity dispersion and variable normalisation or peak value (corresponding to different column densities).
@@
We consider the second moment of the remaining $>4\sigma_{\rm ch}$ profile after masking as a function of the S/N of the profile. Without masking (and for noise-free data), this would equal the velocity dispersion of the Gaussian. With masking, it decreases as the peak value decreases.
@@
Note how well the down-turn in the tail is described by the simple single Gaussian model. We can deduce for the masking procedure described here that at S/N values between $\sim 50$ and $\sim 10$, the second moment values start to underestimate the velocity dispersion. At a ${\rm S/N}\sim 5$, the second moment underestimates the velocity dispersion by a factor of $\sim 2.5$. Column densities, though not as affected, can nevertheless be underestimated by factors of $\sim 1.6$ at these lower S/N values. Multiple components along the line of sight will not significantly influence the lower bounds of the distribution as, at a given column density, these will be found at higher second-moment values. Nevertheless, fitting multiple Gaussians to the velocity profiles could further elucidate (or better: circumvent) the effects of masking in the presence of multiple profiles.
@@ 8 Other galaxies in the MHONGOOSE fields
The sensitivity and large field of view of MeerKAT result in the detection of additional companion and field galaxies. Here we give a first inventory of the galaxies detected in the 1.5 by 1.5 single-track fields in a velocity range of $-500$ to $+500$ km s-1 with respect to the central velocity of the target galaxies. Detections are based on the single-track zeroth-moment maps. A list of these additional galaxies is given in Table 7. For the fields where additional galaxies are present, Fig. 15 shows the positions of these galaxies. We expect an even larger number of galaxies to be detected in the full-depth data once available. While the search volume is defined with respect to our target galaxies, this does not imply that every one of these additional galaxies is also physically associated with those target galaxies. Most of them will be, but it is possible that (especially at the edges of the volume probed) we are picking up some foreground or background galaxies.
@@
Three additional galaxies visible in the moment maps are partially outside of the volume of the data cubes used here. These are LEDA 583148 in the J0309–41 field, ESO 486–21 in the J0459–26 field, and MKT J125225.4-124304.1 in the J1253–12 field. For these three galaxies we list the position and approximate radial velocity in Table 7. We refrain from presenting their H i line widths and fluxes because the large uncertainty due to the incomplete coverage by the data used here. The last of these three galaxies is a fifth, previously uncataloged galaxy. All of the additional galaxies listed here can thus be associated with an optical counterpart; we have not found any “dark” galaxies.
@@
A more complete census of the companion galaxy population will be possible when the full-depth data for the full sample are available and we can probe the companion population to the fullest extent possible for our survey. As with the galaxies presented here, these satellites and companions will be resolved spatially and spectrally, meaning that in-depth studies of the dark matter content and baryon fractions can be made down to an H i mass limit of $\sim 10^{6}$ $M_{\odot}$ (cf. Table 5), a regime previously only accessible in the Local Group. A first discovery of such a low-mass galaxy in the MHONGOOSE observations is presented in Maccagni et al. (2024).
@@
 [^11] Galaxies in bold font match their optical counterparts both in position and velocity. Galaxies in a normal font match the cataloged optical counterpart in position, but with no prior cataloged redshift. For these galaxies, our data give the first redshift measurements. For the galaxies in italic font, there is a spatial match with a previously uncataloged optical counterpart. The “Galaxies with uncertain parameters” are only partially included in the data presented here. The central velocities listed are approximate.
@@ 9 Summary
We present a description of the MHONGOOSE survey, a large, deep, high-resolution H i survey on MeerKAT of 30 nearby gas-rich spiral and dwarf galaxies. The total survey comprises 1650h of observing time, with 55h allocated per galaxy. The main science goal of MHONGOOSE is to characterize the low-column density H i in these galaxies, linking it with various cold-gas-accretion scenarios. Other science goals include detailed explorations of the link between H i and star formation, and the kinematics of the galaxies, constraints on the presence and distribution of dark matter, as well as the environments of the targets. The sample covers approximately three orders of magnitude in H i mass, from low-mass $M_{\rm H{\hskip 0.56917pt\tt I}}\sim 10^{7}$ $M_{\odot}$ as typically found in dwarf galaxies to $M_{\rm H{\hskip 0.56917pt\tt I}}\sim 10^{10}$ $M_{\odot}$ as found in large spiral galaxies. In this paper, we present the survey, the sample, and a summary of the observations. In addition we describe the data reduction procedure and data products.
@@
Each galaxy will eventually be observed for 55h in ten tracks of 5.5h each, and we produce full-depth data cubes at a velocity resolution of 1.4 km s-1 and a range in angular resolutions from $\sim 8^{\prime\prime}$ as the highest resolution to $\sim 90^{\prime\prime}$ as our lowest resolution. The resulting column density sensitivity ($3\sigma$ over 16 km s-1) ranges from $4.3\cdot 10^{19}$ cm-2 at the highest resolution to $4.2\cdot 10^{17}$ cm-2 at the lowest resolution. The $3\sigma$ and 50 km s-1 mass-detection limit is $5.4\cdot 10^{5}$ $M_{\odot}$ at 10 Mpc, which is the median distance of the sample.
@@
We present first results based on the full-depth data of a limited number of galaxies and single-track observations of the complete sample (with the latter observations comprising 10% of the survey data). Comparing full-depth, single-track, and single-dish data, we show that MeerKAT is excellent at recovering the total fluxes of our sample galaxies. We also show that the full-depth data detect on average $\sim 2\%$ more H i than the single-track data. While this increase may seem modest, the additional H i mass it represents is sufficient to allow the existence of a $\sim 10^{18}$ cm-2 low-column-density H i component in the disks of our galaxies. This demonstrates the difficulty in investigating the presence of low-column-density H i using total H i mass measurements.
@@
We also investigated the link between the local H i surface density and the velocity “dispersion” in all of our sample galaxies. We recover the well-known upper limit on H i surface density (measured on kpc scales), and find a lower limit to the measured second-moment value of $\sim 8.5$ km s-1. We find that the H i surface density has two distinct locations in a surface density–velocity dispersion diagram: one at $\sim 5\cdot 10^{20}$ cm-2, representing the gas in the star-forming disks, and one at $\sim 5\cdot 10^{19}$ cm-2 representing gas not participating in star formation. The two regions are distinct and separated at an H i surface density of $2\cdot 10^{20}$ cm-2. Previous studies (e.g., Kanekar et al. 2011) showed that, at this surface density, a rapid change in spin temperature occurs, and that this is the column density where the formation of a cool and dense component of the ISM starts.
@@
We give an overview of all additional galaxies detected in the single-track data. These are defined as the sources detected in H i that were not targeted by the survey but were nevertheless found within the $1.5^{\circ}\times 1.5^{\circ}$ area and spanning 1000 km s-1 in velocity range around the target galaxies. We find that half of these 49 galaxies have cataloged optical counterparts that coincide in position and velocity. Just under half have cataloged optical counterparts that coincide in position but that had until now no previously measured redshifts. Five of our additional galaxies have optical counterparts that have not been previously cataloged. We have not found any H i detections that do not have optical counterparts.
@@
Studies of the full-depth data will be presented in future papers. However, based on the results shown here, it is already clear that the MHONGOOSE data are of excellent quality, showing great promise for future work on the low-column-density H i in nearby galaxies.